# Frequency-Guided Posterior Sampling for Diffusion-Based Image Restoration

### v1

@@ 6.2 Real-World Data
Results on Linear Inverse Problems. Table 1 shows our main results for linear inverse problems. On the Gaussian blurring task, surprisingly even though the forward operator already acts as a low-pass filter, our method significantly improves upon DPS and DSG on the Imagenet dataset. This indicates the benefit of the frequency curriculum on more complex and diverse datasets, a phenomenon also demonstrated qualitatively in Figure 4. On the more complex image restoration tasks, our method outperforms the baselines in almost all metrics. Notably, even though Score-SDE/ILVR uses a sequence of noisy measurements, this sequence is not adapted to the frequency characteristics of the data, which results in lower-quality reconstructions as compared to our method (see Appendix E for further discussion). In Figure 4, we see the qualitative differences between the methods. For high-pass filter operators, existing methods generate visually implausible images, indicating that the approximation errors push the denoised outputs away from the natural image manifold. In contrast, our method is able to accurately reconstruct the underlying ground truth image. For blurry images, we see that DPS tends to produce smoothed images, losing high-frequency details. While DSG is able to improve upon this and generate images with fine-grained characteristics of the underlying image, it tends to also generate artifacts. In contrast, due to our frequency curriculum, our method is able to generate images with high perceptual quality. We give further qualitative comparisons to all baselines in Appendix F.
@@
Results on Image Dehazing. Table 2 shows our main results for the nonlinear image dehazing task on the FFHQ dataset. A key strength of our method is the ability to handle nonlinear inverse problems, unlike methods such as [24, 35, 8]. In Figure 4, we observe that our method successfully reconstructs finer details such as background details and facial characteristics similar to what we observed for linear inverse problems. DSG produces slightly grainy reconstructions, but performs much better on the dehazing task than the other tasks. Due to the powerful generative capacity of diffusion models, our method also performs better than problem-specific supervised methods such as AOD-Net [26], a phenomenon also observed in [33]. We provide more qualitative examples in Appendix F.
@@
Effect of Frequency Curriculum. A key ingredient of our method is the frequency curriculum with a time-varying low pass filter $\mathbf{k}_{t}$. To understand the effect of this curriculum, we compare taking a time-dependent $\mathbf{k}_{t}$ vs. a fixed $\mathbf{k}$ that zeroes out frequencies above $\tau_{1}$ for all $t$. In Figure 5, we see that for the high-pass filter operator, the generated reconstructions are still visually implausible images for a fixed $\mathbf{k}$, indicating the frequency curriculum is crucial to our method. We examine this in detail in Appendix F as well as conduct further ablation studies.
@@ 7 Conclusion
In this paper, we provide a theoretical analysis characterizing the approximation error in existing diffusion-based restoration methods, demonstrating how these errors can become severe when forward operators interact with high-frequency components of images. Then, motivated by these insights, we introduce Frequency Guided Posterior Sampling (FGPS), a method that progressively incorporates frequency information during the restoration process in a data-dependent manner. Our extensive experiments on challenging image restoration tasks, such as motion deblurring and image dehazing, demonstrate that FGPS significantly outperforms existing methods. This work opens up new directions for analyzing and improving diffusion model performance on inverse problems by carefully considering the frequency characteristics of both the data and the degradation process.
@@ A.1 Diffusion Models
[…]
@@
[…]
@@ A.2 Signal Processing
[…]
@@
Appendix B Proofs for Section 4.1
@@
By construction, we have that $\mathcal{F}(\mathbf{R})[f_{k}]=S(f_{k})$. This shows that in expectation, $\mathbf{x}$ follows a power spectral density of $S(f_{k})$. Lastly, because $\bm{\Sigma}$ is circulant, we have that the eigenvalues of $\bm{\Sigma}$ are the Discrete Fourier Transform of the first row, which is $\mathbf{R}$. As before, by construction, we have that $\mathcal{F}(\mathbf{R})[f_{k}]=S(f_{k})$. From Equation 30, we have that the eigenvalues of $\bm{\Sigma}$ are precisely $c|f_{k}|^{-\beta}$. ∎
Before we prove our main result, Theorem 2, we prove a useful lemma that calculates the posterior denoising distribution under a multivariate Gaussian assumption on the data.
@@
$\begin{bmatrix}\mathbf{x}_{0}\\ \mathbf{x}_{t}\end{bmatrix}\sim\mathcal{N}\left(\mathbf{0},\begin{bmatrix}\bm{\Sigma}&\sqrt{\alpha_{t}}\bm{\Sigma}\\ \sqrt{\alpha_{t}}\bm{\Sigma}&\alpha_{t}\bm{\Sigma}+(1-\alpha_{t})\mathbf{I}\end{bmatrix}\right).$ (33)
@@
Next, we prove our main theoretical result, Theorem 2, restated below.
@@
Suppose $\mathbf{x}_{0}$ is drawn from $\mathcal{N}(\mathbf{0},\bm{\Sigma})$ and we are given linear measurements $\mathbf{y}=\mathcal{A}(\mathbf{x}_{0})+\mathbf{z}$, where $\mathcal{A}(\mathbf{x}_{0})=\mathbf{A}\mathbf{x}_{0}$ and $\mathbf{z}\sim\mathcal{N}(0,\sigma_{y}^{2}I)$. Suppose that the intermediate value $\mathbf{x}_{t}$ of the continuous-time reverse diffusion from Equation 5 can be written as $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ where $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Then, we have that
$\!\!\!\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}\mid\mathbf{x}_{t})=(\mathbf{A}\bm{\Gamma}_{t})^{T}\!(\mathbf{A}\bm{\Sigma}_{0\mid t}\mathbf{A}^{T}\!+\sigma_{y}^{2}\bm{I})^{-1}\bm{\Delta}_{t},\!\!$ (36)
where $\bm{\Gamma}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{\Sigma}(\bar{\alpha}_{t}\bm{\Sigma}+(1-\bar{\alpha}_{t})\bm{I})^{-1}$, $\bm{\mu}_{0\mid t}=\mathbb{E}[\mathbf{x}_{0}\mid\mathbf{x}_{t}]=\bm{\Gamma}_{t}\mathbf{x}_{t}$, $\bm{\Sigma}_{0\mid t}=\mathrm{Cov}[\mathbf{x}_{0}\mid\mathbf{x}_{t}]=\bm{\Sigma}-\sqrt{\bar{\alpha}_{t}}\bm{\Gamma}_{t}\bm{\Sigma}$, and $\bm{\Delta}_{t}=\mathbf{y}-\mathbf{A}\bm{\mu}_{0\mid t}$. Moreover, the approximation gap $\mathcal{J}^{\text{DPS}}:=s^{\text{DPS}}(\mathbf{x}_{t},\mathbf{y})-s^{*}(\mathbf{x}_{t},\mathbf{y})$ is given by:
$\mathcal{J}^{\text{DPS}}=\left(\mathbf{A}\bm{\Gamma}_{t}\right)^{T}(\sigma_{y}^{-2}\mathbf{I}-(\mathbf{A}\bm{\Sigma}_{0\mid t}\mathbf{A}^{T}+\sigma_{y}^{2}\mathbf{I})^{-1})\bm{\Delta}_{t}.$ (37)
@@
$\begin{bmatrix}\mathbf{x}_{0}\\ \mathbf{y}\end{bmatrix}\Big{|}\overleftarrow{\mathbf{x}_{t}}\sim\mathcal{N}\left(\begin{bmatrix}\bm{\mu}_{0\mid t}\\ \mathbf{A}\bm{\mu}_{0\mid t}\end{bmatrix},\begin{bmatrix}\bm{\Sigma}_{0\mid t}&\bm{\Sigma}_{0\mid t}\mathbf{A}^{T}\\ \mathbf{A}\bm{\Sigma}_{0\mid t}&\mathbf{A}\bm{\Sigma}_{0\mid t}\mathbf{A}^{T}+\sigma_{y}^{2}\mathbf{I}\end{bmatrix}\right).$ (38)
@@
As $\bm{\mu}_{0\mid t}=\bm{\Gamma}_{t}\overleftarrow{\mathbf{x}_{t}}$, we have that $\frac{\partial\bm{\mu}_{0\mid t}}{\partial\overleftarrow{\mathbf{x}_{t}}}=\bm{\Gamma}_{t}$, which gives us Equation 36. The DPS approximation to the true conditional score is $\nabla_{\overleftarrow{\mathbf{x}_{t}}}\log p(\mathbf{y}\mid\overleftarrow{\mathbf{x}_{t}})\approx\nabla_{\overleftarrow{\mathbf{x}_{t}}}\log p(\mathbf{y}\mid\bm{\mu}_{0\mid t})=\nabla_{\overleftarrow{\mathbf{x}_{t}}}\mathcal{N}(\mathbf{A}\bm{\mu}_{0\mid t},\sigma_{y}^{2}\mathbf{I})$. Similar to above, we can also calculate its gradient with respect to $\overleftarrow{\mathbf{x}_{t}}$ as
$\!\!\ \nabla_{\overleftarrow{\mathbf{x}_{t}}}\log p(\mathbf{y}\mid\bm{\mu}_{0\mid t})=\left(\mathbf{A}\bm{\Gamma}_{t}\right)^{T}(\sigma_{y}^{-2}\mathbf{I})\bm{\Delta}_{t}$ (42)
The approximation gap $\mathcal{J}^{\text{DPS}}$ is defined as
$\mathcal{J}^{\text{DPS}}$ $=s^{\text{DPS}}(\mathbf{x}_{t},\mathbf{y})-s^{*}(\mathbf{x}_{t},\mathbf{y})$ (43)
$=\nabla_{\overleftarrow{\mathbf{x}_{t}}}\log p(\mathbf{y}\mid\bm{\mu}_{0\mid t})-\nabla_{\overleftarrow{\mathbf{x}_{t}}}\log p(\mathbf{y}\mid\mathbf{x}_{t})$ (44)
Note that in the second line, the unconditional score term cancels, so we are left only with the noisy likelihood difference. Subtracting Equation 36 from Equation 42, we have the approximation gap $\mathcal{J}^{\text{DPS}}$ is equal to $\left(\mathbf{A}\bm{\Gamma}_{t}\right)^{T}(\sigma_{y}^{-2}\mathbf{I}-(\mathbf{A}\bm{\Sigma}_{0\mid t}\mathbf{A}^{T}+\sigma_{y}^{2}\mathbf{I})^{-1})\bm{\Delta}_{t}$, completing our proof. ∎
@@
Appendix C Further Motivations for our Method
In Section 4.1, we argued that when the forward operator is convolution with a high-pass filter, existing methods have a large approximation between the conditional score and its approximation, which can lead to compromised sample quality in practice. High-pass filtering may seem like a contrived example, because after all, many inverse problem tasks considered in the literature have forward operators that are low-pass filters, such as Gaussian deblurring and superresolution tasks. However, for complex image restoration tasks such as motion deblurring, the forward operator can indeed act as a high-pass filter in certain spatial directions of the image. In Figure 6, we examine this effect by looking at the log magnitude of the frequency domain of an image convolved with a simple directional blur kernel. We can clearly see that in the red circled direction of the Fourier domain, the filter retains high frequency components of the original image. This is also mathematically evident using the Fourier convolution theorem. The Fourier convolution theorem states that

### v2

@@ 6.2 Results
Results on Linear Inverse Problems. Table˜1 shows our main results for linear inverse problems. On the FFHQ dataset, for simple tasks such as Gaussian deblurring, FGPS gives comparable performance to baselines, which already have strong reconstruction performance. However, on more complex image restoration tasks such as motion deblurring, our method outperforms the baselines in almost all metrics. Notably, even though Score-SDE/ILVR uses a sequence of noisy measurements, this sequence is not adapted to the frequency characteristics of the data, which results in lower quality reconstructions as compared to our method (see Appendix F for further discussion). In Figure˜3, we see the qualitative differences between the methods. For high-pass filter operators, existing methods generate visually implausible images, indicating that the approximation errors push the denoised outputs away from the natural image manifold. In contrast, our method is able to accurately reconstruct the underlying ground truth image. For blurry images, we see that DPS tends to produce smoothed images, losing high-frequency details. While DSG is able to improve upon this and generate images with fine-grained characteristics of the underlying image, it tends to also generate artifacts. In contrast, due to our frequency curriculum, our method is able to generate images with high perceptual quality. We give further qualitative comparisons to baselines in Appendix G.
Role of Dataset. For the ImageNet dataset, we observe that FGPS significantly outperforms all baselines, especially in perceptual quality. We conjecture that this is because on more complex datasets with varying frequency characteristics, the explicit frequency schedule for the measurement is essential to ensure the reverse process iterates do not deviate far from the intermediate noisy data manifolds learned by the diffusion models.
@@
Results on Image Dehazing. Table 2 shows our main results for the nonlinear image dehazing task on the FFHQ dataset. A key strength of our method is the ability to handle nonlinear inverse problems, unlike methods such as [24, 37, 8]. In Figure˜3, we observe that our method successfully reconstructs finer details such as background details and facial characteristics similar to what we observed for linear inverse problems. DSG produces slightly grainy reconstructions, but performs much better on the dehazing task than the other tasks. Due to the powerful generative capacity of diffusion models, our method also performs better than problem-specific supervised methods such as AOD-Net [26], a phenomenon also observed in [35]. We provide more qualitative examples in Appendix G.
@@
Figure 4: Visualization of Transformed Measurements $y_{t}$ at different timesteps $t$, demonstrating our coarse-to-fine strategy.
Visualizing Measurement Sequence. In Figure˜4, we show the transformed measurements at three timesteps in the reverse diffusion process when running FGPS. In the early stages of the reverse process, the measurement retains only a coarse outline of the measurement. This coarse-to-fine strategy is essential to ensuring $\mathbf{x}_{t}$ remains on the data manifold learned by the diffusion model.
Effect of Frequency Curriculum. A key ingredient of our method is the frequency curriculum with a time-varying low pass filter $\mathbf{k}_{t}$. To understand the effect of this curriculum, we compare taking a time-dependent $\mathbf{k}_{t}$ vs. a fixed $\mathbf{k}$ that zeroes out frequencies above $\tau_{1}$ for all $t$. In Figure˜5, we see that for the high-pass filter operator, the generated reconstructions are still visually implausible images for a fixed $\mathbf{k}$, indicating the frequency curriculum is crucial to our method. Our method only requires $\mathbf{k}_{1}$ to be a low frequency cut-off, otherwise the reconstructions look like Figure˜5. This corroborates our theoretical results, which show that the approximation gap early on in the reverse process greatly influences reconstruction quality. Finally, our method does not qualitatively change with higher cutoffs for $\mathbf{k}_{T}$. We examine this in detail in Appendix G as well as conduct further ablation studies.
@@ 7 Conclusion
In this paper, we introduce Frequency Guided Posterior Sampling (FGPS), a method that progressively incorporates frequency information in a data-dependent manner. We provide a theoretical analysis characterizing the approximation error of FGPS and existing diffusion-based restoration methods, demonstrating how these errors can become severe when forward operators interact with high-frequency components of images. Our experiments on challenging image restoration tasks, such as motion deblurring and image dehazing, demonstrate that FGPS significantly outperforms existing methods. This work opens up new directions for analyzing diffusion-based inverse problem solvers by carefully considering the frequency characteristics of both the data and the degradation process.
Acknowledgements
D.T. was supported by a gift from AWS to Penn Engineering’s ASSET Center for Trustworthy AI. R.V. thanks the support of Penn startup funds and the Research Collaboration on the Mathematical and Scientific Foundations of Deep Learning under grants NSF 2031985 and Simons 814201.
@@ A.1 Diffusion Models
$=\frac{1}{p(\mathbf{x}_{t})}\int\nabla_{\mathbf{x}_{t}}p(\mathbf{x}_{t}\mid\mathbf{x}_{0})p(\mathbf{x}_{0})\ d\mathbf{x}_{0}.$ (18)
@@
$=\int p(\mathbf{x}_{0}\mid\mathbf{x}_{t})\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}\mid\mathbf{x}_{0})\ d\mathbf{x}_{0}$ (19)
$=\int p(\mathbf{x}_{0}\mid\mathbf{x}_{t})\frac{\sqrt{\alpha_{t}}\mathbf{x}_{0}-\mathbf{x}_{t}}{1-\alpha_{t}}\ d\mathbf{x}_{0}$ (20)
@@ A.2 Signal Processing
$\mathcal{F}_{\text{cont}}(\mathbf{x})[f]=\int_{-\infty}^{\infty}\mathbf{x}(t)e^{-2\pi ift}dt,$ (22)
@@
Appendix B Proofs for Section˜5
@@
By construction, we have that $\mathcal{F}(\mathbf{R})[f_{k}]=S(f_{k})$. This shows that in expectation, $\mathbf{x}$ follows a power spectral density of $S(f_{k})$. Lastly, because $\bm{\Sigma}$ is circulant, we have that the eigenvalues of $\bm{\Sigma}$ are the Discrete Fourier Transform of the first row, which is $\mathbf{R}$. As before, by construction, we have that $\mathcal{F}(\mathbf{R})[f_{k}]=S(f_{k})$. From Equation˜29, we have that the eigenvalues of $\bm{\Sigma}$ are precisely $c|f_{k}|^{-\beta}$. ∎
Before we prove our main result, Theorem˜2, we prove a useful lemma that calculates the posterior denoising distribution under a multivariate Gaussian assumption on the data.
@@
[…]
@@
Next, we prove our main theoretical result, Theorem˜2, restated below.
@@
Suppose $\mathbf{x}_{0}$ is drawn from $\mathcal{N}(\mathbf{0},\bm{\Sigma})$ and we are given linear measurements $\mathbf{y}=\mathcal{A}(\mathbf{x}_{0})+\mathbf{z}$, where $\mathcal{A}(\mathbf{x}_{0})=\mathbf{A}\mathbf{x}_{0}$ and $\mathbf{z}\sim\mathcal{N}(0,\sigma_{y}^{2}I)$. Suppose that the intermediate value $\mathbf{x}_{t}$ of the continuous-time reverse diffusion from Equation˜5 can be written as $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ where $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Then, we have that the true noisy likelihood score $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}\mid\mathbf{x}_{t})$ is
$(\mathbf{A}\bm{\Gamma}_{t})^{T}\!(\mathbf{A}\bm{\Sigma}_{0\mid t}\mathbf{A}^{T}\!+\sigma_{y}^{2}\bm{I})^{-1}(\mathbf{y}-\mathbf{A}\bm{\mu}_{0\mid t}),\!\!$ (35)
where $\bm{\Gamma}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{\Sigma}(\bar{\alpha}_{t}\bm{\Sigma}+(1-\bar{\alpha}_{t})\bm{I})^{-1}$, $\bm{\mu}_{0\mid t}=\mathbb{E}[\mathbf{x}_{0}\mid\mathbf{x}_{t}]=\bm{\Gamma}_{t}\mathbf{x}_{t}$ and $\bm{\Sigma}_{0\mid t}=\mathrm{Cov}[\mathbf{x}_{0}\mid\mathbf{x}_{t}]=\bm{\Sigma}-\sqrt{\bar{\alpha}_{t}}\bm{\Gamma}_{t}\bm{\Sigma}$. Moreover, the FGPS approximation $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{C}_{t}\mathbf{y}\mid\bm{\mu}_{0\mid t})$ can be analytically calculated as
$\left(\mathbf{C}_{t}\mathbf{A}\bm{\Gamma}_{t}\right)^{T}(\sigma_{y}^{2}\mathbf{C}_{t}\mathbf{C}_{t}^{T})^{-1}(\mathbf{C}_{t}\mathbf{y}-\mathbf{C}_{t}\mathbf{A}\bm{\mu}_{0\mid t}).$ (36)
@@
[…]
@@
As $\bm{\mu}_{0\mid t}=\bm{\Gamma}_{t}\overleftarrow{\mathbf{x}_{t}}$, we have that $\frac{\partial\bm{\mu}_{0\mid t}}{\partial\overleftarrow{\mathbf{x}_{t}}}=\bm{\Gamma}_{t}$, which gives us Equation˜35. The FGPS approximation to the true conditional score is $\nabla_{\overleftarrow{\mathbf{x}_{t}}}\log p(\mathbf{y}\mid\overleftarrow{\mathbf{x}_{t}})\approx\nabla_{\overleftarrow{\mathbf{x}_{t}}}\log p(\mathbf{C}_{t}\mathbf{y}\mid\bm{\mu}_{0\mid t})=\nabla_{\overleftarrow{\mathbf{x}_{t}}}\mathcal{N}(\mathbf{A}\bm{\mu}_{0\mid t},\sigma_{y}^{2}\mathbf{I})$. Similar to above, we can also calculate its gradient $\nabla_{\overleftarrow{\mathbf{x}_{t}}}\log p(\mathbf{C}_{t}\mathbf{y}\mid\bm{\mu}_{0\mid t})$ with respect to $\overleftarrow{\mathbf{x}_{t}}$ as
[…]
Corollary 3 can easily be proven by taking $\mathbf{C}_{t}$ as the identity matrix in the above proof.
@@
Appendix C Theoretical Investigation of Approximation Gap of FGPS and DPS
Our main theorem shows in the multivariate Gaussian setting, the true conditional score differs from approximations in the term $(\mathbf{I}+\mathbf{A}\bm{\Sigma}_{0\mid t}\mathbf{A}^{\top})^{-1}$, which requires approximations in general. DPS approximates it at as the identity matrix. In this section, we show that in certain cases, FGPS is a significantly better approximation than the identity matrix. To see this, consider when $\mathbf{A}$ is a high-pass filter and the data covariance is $\bm{\Sigma}_{f}$ such that it follows a power law in the frequency domain. Then, the matrix $\mathbf{A}\bm{\Sigma}_{0\mid t}\mathbf{A}^{\top}$ has significant energy in high-frequency directions. Specifically, in the Fourier basis, denoting the eigenvalues of $\mathbf{A}$, $\bm{\Sigma}_{f}$, $\bm{\Sigma}_{0\mid t}$ as $a_{i}$, $\lambda_{i}$ and $\lambda_{i}^{t}$ respectively, the eigenvalues of $\mathbf{A}\bm{\Sigma}_{0\mid t}\mathbf{A}^{\top}$ take the form $a_{i}^{2}\lambda_{i}^{t}$, where $a_{i}^{2}$ grows with frequency (due to the high-pass nature of $\mathbf{A}$), and $\lambda_{i}^{t}=\frac{(1-\alpha_{t})\lambda_{i}}{\alpha_{t}\lambda_{i}+(1-\alpha_{t})}$. Although $\lambda_{i}^{t}$ is small in high-frequency directions, the product $a_{i}^{2}\lambda_{i}^{\text{post}}$ can still be $\mathcal{O}(1)$ or larger due to the amplification by $\mathbf{A}$. Consequently, the eigenvalues of $(\mathbf{I}+\mathbf{A}\bm{\Sigma}_{0\mid t}\mathbf{A}^{\top})^{-1}$ in those directions, given by $\frac{1}{1+a_{i}^{2}\lambda_{i}^{\text{post}}}$, are significantly less than $1$, indicating strong suppression of high-frequency components especially when $\alpha_{t}$ is very small as in the beginning of the reverse process. In contrast, low-frequency directions (where $a_{i}^{2}\approx 0$) are preserved. This shows that $(\mathbf{I}+\mathbf{A}\bm{\Sigma}_{0\mid t}\mathbf{A}^{\top})^{-1}$ acts as a low-pass filter. The FGPS approximation, $\mathbf{C}_{t}^{\top}(\mathbf{C}_{t}\mathbf{C}_{t}^{\top})^{-1}\mathbf{C}_{t}$, which is a projection matrix onto the low-frequency components, thus approximates this behavior more faithfully than the identity matrix, which uniformly preserves all directions.
Appendix D Further Motivations for our Method
In Section˜5, we argued that when the forward operator is convolution with a high-pass filter, existing methods have a large approximation between the conditional score and its approximation, which can lead to compromised sample quality in practice. High-pass filtering may seem like a contrived example, because after all, many inverse problem tasks considered in the literature have forward operators that are low-pass filters, such as Gaussian deblurring and superresolution tasks. Even though some natural imaging systems such as phase contrast microscopy, we argue that the high-pass filter effect can also show up in more complex image restoration tasks. For example, for motion deblurring, the forward operator can act as a high-pass filter in certain spatial directions of the image. In Figure˜6, we examine this effect by looking at the log magnitude of the frequency domain of an image convolved with a simple directional blur kernel. We can clearly see that in the red circled direction of the Fourier domain, the filter retains high frequency components of the original image. This is also mathematically evident using the Fourier convolution theorem. The Fourier convolution theorem states that